The rms speed of the molecules in 1.7 g of hydrogen gas is 1800 m/s.
A)What is the total translational kinetic energy of the gas molecules?
B)What is the thermal energy of the gas?
C)500 J of work are done to compress the gas while, in the same process, 1800 J of heat energy are transferred from the gas to the environment. Afterward, what is therms speed of the molecules?
Answer #1

(a) translational kinetic energy = 1/2mv^2 = 1/2(1.7/1000)(1800)^2 = 2754 J
(b) thermal energy = 5/3 * 1/2mv^2 = 4590 kJ

(c)

total theral energy remaining:

= 4590 +500-2000 = 3090 J

Translaiona energy:

= 3/5*3090 =1854 J

Hence,

1854 =1/2mv^2 = 1/2(1.7/1000)(v)^2

v = 1476.88 m/s